Please help.
Solve each of the following systems by addition. If a unique solution does not exist, state whether the system is inconsistent or dependent.
2x + 3y = 1
5x + 3y = 16
Thanks
JMS

Answer by Adam(64)

Multiply the second equation by -1
and get
2x + 3y = 1
-5x - 3y = -16
add second equation to first
-3x + 0y = -15
-3x = -15
3x = 15
x = 5
substitute this into any of the original equations
2*5 + 3y = 1
3y = -9
y = -3